# Teaching Code LLMs to Use Autocompletion Tools in Repository-Level Code Generation

### v2

@@ V-A RQ1: Similarity Metrics
SUMMARY: Our approach, ToolGen, consistently achieves significantly higher Static Validity Rate in repository-level lint analysis compared to the baselines, with improvements ranging from 10.9% to 42.2%. These results underscore the effectiveness of our tool-integrated generation process in mitigating the generation of invalid identifiers, a common challenge faced by conventional code LLMs in the context of repository-level code generation.
V-D RQ3: Generalizability
@@
Based on the results presented in Table II and Table III, our tool-integrated generation approach consistently improves performance in repository-level metrics while maintaining comparative similarity metrics across various model architectures (decoder-only and encoder-decoder) and parameter scales (ranging from 124 million to 7 billion).
Specifically, for the encoder-decoder model CodeT5, our tool-integrated generation approach demonstrates well-balanced and comprehensive enhancements in repository-level metrics. However, for the two decoder-only models, CodeGPT and CodeLlama, the improvements achieved by ToolGen vary between Dependency Coverage and Static Validity Rate. When employing the smaller CodeGPT as the base model, ToolGen significantly enhances Dependency Coverage but exhibits less pronounced improvements in the Static Validity Rate. In contrast, when employing the larger CodeLlama, the results are reversed. Upon closer examination of the generated functions, we attribute these variations to CodeGPT’s occasional difficulty in predicting the <EOS> token correctly, leading to syntax errors that violate the prerequisites for successful lint check. In the case of CodeLlama, the initially high Dependency Coverage, owing to its extensive parameter scale and training data, poses challenges for improvement through parameter-efficient fine-tuning (PEFT) with LoRA (only 0.25% trainable parameters).
@@
SUMMARY: Our tool-integrated generation approach consistently improves repository-level metrics while maintaining competitive similarity metrics across various model architectures and parameter scales. This suggests that our approach is versatile and has the potential for broader applicability with other base models in repository-level code generation.
@@ VI Discussion
VI-A Limitations
@@
Repository-level Evaluation Metrics: In the computation of the two repository-level evaluation metrics, namely dependency coverage and success rate, we employ static analysis tools to identify target expressions and perform lint examinations. Similar to the autocompletion tools, these static tools may introduce a degree of inaccuracy into the calculated metrics. However, it is essential to note that this does not significantly impact the demonstrated effectiveness of ToolGen, as the baseline metrics are also determined using the same static tools. Another limitation is that our evaluation does not encompass metrics such as pass rate or pass@k for test cases, which are commonly used to assess the logical correctness of standalone functions generated by code LLMs. The primary reason for this omission is that such metrics typically rely on the execution of manually crafted unit test cases, which can be challenging to carry out in the context of repository-level code generation. Instead, our focus in this work centers on mitigating dependency errors, such as no-member and undefined-variable errors, stemming from the unawareness of repository-level dependencies.
VI-B Threats to Validity
Internal Threats: The first internal threat pertains to potential data quality issues common in learning-based approaches. To mitigate this threat, we construct our augmented dataset and evaluation benchmark dataset using the widely adopted CodeSearchNet dataset, which serves as a reliable source for pretraining and evaluating various code models. Another internal threat pertains to the potential data leakage for CodeLlama, as the code repositories in the benchmark dataset may have been encountered by CodeLlama during its pretraining phase. However, our generalizability evaluation (RQ3) provides evidence of consistent performance across the three ToolGen variants, suggesting that the improvements achieved by ToolGen-llama in repository-level code generation are not attributed to data leakage.
@@ VII-B Incorporating External Tools into LLMs
Recent research [20, 21, 36, 37, 38, 39, 40, 41, 42, 43, 44] has explored the integration of external tools (e.g., search engines, web browsers, calculators, and python interpreters) into the LLM generation process, aiming to address their limitations in certain generation scenarios. For instance, Schick et al. propose ToolFormer [20], which augments datasets to instruct LLMs on invoking existing arithmetic calculators, effectively reducing errors in generated text related to arithmetic calculations. Building upon this idea, Zhang et al. introduce ToolCoder [21], designed to teach LLMs how to utilize information-retrieval-based (IR-based) API search tools during code generation. While ToolCoder is effective in generating functionally correct standalone functions, it falls short in addressing repository-level dependencies, limiting its ability to resolve dependency errors. Another relevant example is Repilot [22], which utilizes code completion tools to filter out impractical suggestions made by LLMs in the context of automatic program repairing (APR), focusing on generating valid bug-fix patches rather than entire functions.
@@ VIII Conclusion
We present ToolGen, a novel approach that seamlessly integrates autocompletion tools into the code LLM generation process to effectively address repository-level dependencies. ToolGen encompasses two crucial phases: Data Augmentation and Model Fine-tuning, and Tool-integrated Code Generation. Our comprehensive evaluation showcases ToolGen’s substantial improvements in the two introduced repository-level metrics across three distinct code LLMs, while also demonstrating its competitiveness in widely-recognized similarity metrics. Moreover, our generalizability evaluation reaffirms ToolGen’s consistent performance when applied to diverse code LLMs, including various model architectures and scales.

### v3

@@ V-B2 Static Validity Rate
SUMMARY: Our approach, ToolGen, consistently achieves significantly higher Static Validity Rate in repository-level lint analysis compared to the baselines, with improvements ranging from 44.9% to 57.7%. These results underscore the effectiveness of our tool-integrated generation process in mitigating the generation of invalid identifiers, a common challenge faced by conventional code LLMs in the context of repository-level code generation.
V-C RQ3: Execution-based Effectiveness
Table IV presents the detailed evaluation results for test case execution (Pass@1) on the 176 CoderEval coding tasks.
TABLE IV: Evaluation results of execution-based effectiveness. self, slib, plib, class, file, and project represent self-contained, slib-runnable, plib-runnable, class-runnable, file-runnable, and project-runnable, respectively. The numbers in brackets after each runnable level indicate the corresponding number of tasks, while the numbers in brackets after the rates indicate the number of generated functions that pass the test cases. In each base model group, the best results are highlighted in gray, except when all results are the same.
  Approach Execution-based Metric (Pass@1)
total (176) self (26) slib (23) plib (15) class (49) file (51) project (12)
  Vanilla-gpt 3.4% (6) 7.7% (2) 4.3% (1) 6.7% (1) 2.0% (1) 2.0% (1) 0.0% (0)
ToolGen-gpt (ours) 3.4% (6) 3.8% (1) 4.3% (1) 6.7% (1) 2.0% (1) 3.9% (2) 0.0% (0)
[…]
ragToolGen-gpt (ours) 2.8% (5) 0.0% (0) 4.3% (1) 6.7% (1) 4.1% (2) 2.0% (1) 0.0% (0)
Vanilla-t5 4.0% (7) 7.7% (2) 4.3% (1) 13.3% (2) 4.1% (2) 0.0% (0) 0.0% (0)
ToolGen-t5 (ours) 5.1% (9) 15.4% (4) 8.7% (2) 6.7% (1) 4.1% (2) 0.0% (0) 0.0% (0)
[…]
ragToolGen-t5 (ours) 5.1% (9) 15.4% (4) 8.7% (2) 6.7% (1) 4.1% (2) 0.0% (0) 0.0% (0)
Vanilla-llama 6.8% (12) 19.2% (5) 13.0% (3) 13.3% (2) 4.1% (2) 0.0% (0) 0.0% (0)
ToolGen-llama (ours) 8.5% (15) 23.1% (6) 13.0% (3) 20.0% (3) 4.1% (2) 2.0% (1) 0.0% (0)
[…]
ragToolGen-llama (ours) 10.8% (19) 34.6% (9) 8.7% (2) 6.7% (1) 6.1% (3) 0.0% (0) 33.3% (4)
Comparison to Vanilla baselines. Compared to the three Vanilla baselines, our approach ToolGen generates 0, 2, and 3 additional functionally-correct functions, resulting in 0%, 40.0%, and 25.0% improvements in Pass@1, respectively. Specifically, ToolGen-gpt improves the pass rate for file-runnable tasks, while reducing pass rate for self-contained tasks; ToolGen-t5 improves pass rate for slib-runnable file-runnable tasks; ToolGen-llama improves pass rate for self-contained, plib-runnable, and class-runnable tasks. These tasks require different runnable-level dependencies (such as local variables and user-defined functions) to achieve correct functionality in the code. The enhancements by ToolGen underscore the effectiveness of integrating autocompletion tools to handle these dependencies.
Comparison to RepoCoder baselines. Overall, the RepoCoder baselines show unstable performance across different base models. Specifically, compared to their respective Vanilla baselines, RepoCoder-gpt and RepoCoder-t5 exhibit reductions or no improvement in test pass rates, while RepoCoder-llama shows significant improvement. When compared to RepoCoder-gpt and RepoCoder-t5, our ToolGen-gpt and ToolGen-t5 show improvements in pass rates with 1 and 2 more functions passing the test cases, respectively. However, ToolGen-llama exhibits a lower pass rate than RepoCoder-llama (15 vs. 19). These variations can be attributed to several factors: CodeGPT and CodeT5 have fewer parameters (124 million and 220 million) and stricter token number limitations (1,024 and 512), limiting their ability to process and understand retrieval-augmented prompts. In contrast, CodeLlama, with more model parameters (7 billion) and support for longer token sequences (16,384 tokens), allows RepoCoder-llama to achieve a higher pass rate than Vanilla-llama and ToolGen-llama due to the benefits of RAG.
Integration with RAG. When integrating ToolGen with RAG, ragToolGen-gpt and ragToolGen-t5 do not show improvement, while ragToolGen-llama exhibits breakthroughs for project-runnable tasks, with 4 more generated functions passing the test cases. However, the overall pass rate remains unchanged after integrating RAG, showing different advantages and disadvantages of RAG integration for various runnable-level dependencies.
SUMMARY: Our approach, ToolGen, outperforms or matches the three Vanilla baselines (Vanilla-gpt, Vanilla-t5, and Vanilla-llama) by generating 0, 2, and 3 more functionally correct functions, achieving 0%, 40.0%, and 25.0% improvements in Pass@1, respectively. Compared to RAG-based RepoCoder baselines, both ToolGen and RepoCoder have their own advantages and disadvantages for different base models and runnable-level dependencies. Additionally, our decoding-stage tool integration approach shows potential when combined with prompt-level RAG techniques for addressing certain types of dependencies.
V-D RQ4: Efficiency
Figure 6 illustrates the efficiency evaluation conducted on a single NVIDIA H100 Tensor Core GPU (80GB GPU Memory). Our approach exhibits approximately twice the average generation time for the 176 tasks in the CoderEval dataset, ranging from 0.63 to 2.34 seconds across different base models.
Figure 6: Results of efficiency evaluation.
The high efficiency of our tool integration is attributed to the offline trigger insertion and fine-tuning. The autocompletion tool is triggered only when the fine-tuned models predict the trigger token <COMP>, significantly reducing unnecessary tool invocations. Specifically, the fine-tuned CodeGPT, CodeT5, and CodeLlama predict an average of 5.02, 6.24, and 7.05 <COMP> tokens per task, respectively, which is much fewer than the average function length. Additionally, during the generation of a function, autocompletion is often triggered multiple times for the same objects (e.g.,self”); we maintain a cache to recall completion suggestions for previously visited objects, thereby avoiding repeated tool invocations for the same objects.
SUMMARY: Our tool-integrated generation approach, ToolGen, demonstrates high efficiency in repository-level code generation, with latency ranging from 0.63 to 2.34 seconds for generating each function. This efficiency is attributed to predicting the trigger token <COMP> and implementing a caching mechanism for completion suggestions.
V-E RQ5: Generalizability
@@
Based on the results presented in Table II and Table III, our tool-integrated generation approach consistently enhances performance in dependency-based metrics while maintaining comparable similarity-based metrics across various model architectures (decoder-only and encoder-decoder) and parameter scales (ranging from 124 million to 7 billion). According to Table IV and Figure 6, our approach improves or maintains execution-based metrics across the base models, with a consistent and acceptable additional latency overhead.
@@
SUMMARY: Our tool-integrated generation approach consistently improves or maintains dependency-based and execution-based metrics while achieving competitive similarity-based metrics across various model architectures and parameter scales. This suggests that our approach is versatile and has the potential for broader applicability with other base models in repository-level code generation.
@@ VI Discussion
VI-A Case Study
Figure 7 depicts three specific examples using Vanilla-llama and ToolGen-llama. Each row corresponds to an example, presenting the description, ground truth, code generated by Vanilla-llama, and code generated by ToolGen-llama.
Figure 7: Case study of three specific examples. Additional explanatory notes are marked with gray boxes. In the notes, "dep: xxx" denotes a dependency necessary in the generated code.
Example 1: The code generated by ToolGen-llama successfully predicts the member “_value” in the class “Counter”, while Vanilla-llama incorrectly predicts an undefined member “value”, resulting in a no-member error. This difference can be attributed to ToolGen’s integration of the autocompletion tool, which helps the code LLMs recognize necessary dependencies like user-defined attributes/members.
[…]
Example 3: There are two crucial dependencies, namely “cls._get_service()” and “ServiceName.PLUGINS_MANAGER”, necessary to realize the required functionality. Vanilla-llama fails to predict both dependencies and instead outputs non-existent dependencies like “cls._plugins_manager” and “PluginManager()”, causing the generated code to fail lint checks and test cases. For ToolGen-llama, although it successfully predicts the dependency “cls._get_service()”, it fails to predict “ServiceName.PLUGINS_MANAGER” because the model chooses “cls” instead of “ServiceName” when starting predicting the argument for “cls._get_service()”. This misleads the generation in an incorrect direction, resulting in the failure of the final code. This example also highlights the challenges of applying code LLMs in practical code generation, even when integrating autocompletion tools to avoid certain dependency issues. Introducing an incorrect token at any critical step in the generation process can result in the production of erroneous code.
VI-B Limitations
@@
Dependency-based Evaluation Metrics: In the computation of the two repository-level evaluation metrics, namely Dependency Coverage and Static Validity Rate, we employ static analysis to identify target expressions and perform lint examinations. Similar to the autocompletion tools, these static tools may introduce a degree of inaccuracy into the calculated metrics. However, it is essential to note that this does not significantly impact the demonstrated effectiveness of ToolGen, as the baseline metrics are also determined using the same static analysis.
Integration and Comparison with SOTA Closed-source LLMs: Our approach can be applied to any encoder-decoder or decoder-only models. However, for the most state-of-the-art (SOTA) LLMs like GPT-3.5 and GPT-4, integrating the trigger insertion and tool-integrated decoding process faces challenges due to their closed-source nature and impracticality for fine-tuning. In the future, we may explore the possibility of integrating autocompletion tools into such closed-source LLMs in a non-tuning manner. In our evaluation, we do not compare ToolGen with these SOTA closed-source LLMs, as our goal is to assess the effectiveness of integrating autocompletion tools for repository-level code generation. Therefore, we focus on comparing the performance of ToolGen, Vanilla, and RepoCoder under the same base models.
VI-C Threats to Validity
Internal Threats: The first internal threat pertains to potential data quality issues common in learning-based approaches. To mitigate this threat, we construct our augmented dataset and evaluation benchmark dataset using the widely adopted CodeSearchNet dataset, which serves as a reliable source for pretraining and evaluating various code models. Another internal threat pertains to the potential data leakage for CodeLlama, as the code repositories in the benchmark dataset may have been encountered by CodeLlama during its pretraining phase. However, our generalizability evaluation (RQ5) provides evidence of consistent performance across the three ToolGen variants, suggesting that the improvements achieved by ToolGen-llama in repository-level code generation are not attributed to data leakage.
@@ VII-B Incorporating External Tools into LLMs
Recent research [30, 31, 50, 51, 52, 53, 54, 55, 56, 57, 58] has explored the integration of external tools (e.g., search engines, web browsers, calculators, and python interpreters) into the LLM generation process, aiming to address their limitations in certain generation scenarios. For instance, Schick et al. propose ToolFormer [30], which augments datasets to instruct LLMs on invoking existing arithmetic calculators, effectively reducing errors in generated text related to arithmetic calculations. Building upon this idea, Zhang et al. introduce ToolCoder [31], designed to teach LLMs how to utilize information-retrieval-based (IR-based) API search tools during code generation. While ToolCoder is effective in generating functionally correct standalone functions, it falls short in addressing repository-level dependencies, limiting its ability to resolve dependency errors. More relevant examples are Repilot [41], STALL+ [59], and MGD [60], which utilize code completion tools to filter out impractical suggestions made by LLMs, focusing on generating API/line-level code completions and valid bug-fix patches rather than entire functions.
@@ VIII Conclusion
We present ToolGen, a novel approach that seamlessly integrates autocompletion tools into the code LLM generation process to effectively address repository-level dependencies. ToolGen encompasses two crucial phases: Data Augmentation and Model Fine-tuning, and Tool-integrated Code Generation. Our comprehensive evaluation showcases ToolGen’s improvements in the two introduced dependency-level metrics and a widely used execution-based metric across three distinct code LLMs, while also demonstrating its competitiveness in widely-recognized similarity metrics. ToolGen also demonstrates high efficiency in repository-level code generation, due to the offline fine-tuning with trigger insertion. Moreover, our generalizability evaluation reaffirms ToolGen’s consistent performance when applied to diverse code LLMs, including various model architectures and scales.